OpenStudy (anonymous):

What is the equation, in slope-intercept form, of the line that passes through (0, −2) and has a slope of −3?

OpenStudy (anonymous):

pls help me

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1)\] m=slope (x1,y1)

OpenStudy (anonymous):

is it y = -3x-2?

pooja195 (pooja195):

y+2=-3(x-0) y+2=-3x-0 y=-3x-2

pooja195 (pooja195):

yep thats correct

OpenStudy (anonymous):

ok thank you!

pooja195 (pooja195):

:)